# ECLeKTic: a Novel Challenge Set for Evaluation of Cross-Lingual Knowledge Transfer

### v2

@@ Abstract
To achieve equitable performance across languages, multilingual large language models (LLMs) must be able to abstract knowledge beyond the language in which it was acquired. However, the current literature lacks reliable ways to measure LLMs’ capability of cross-lingual knowledge transfer. To that end, we present ECLeKTic,111The dataset is available at https://www.kaggle.com/datasets/googleai/eclektic a multilingual closed-book QA (CBQA) dataset that Evaluates Cross-Lingual Knowledge Transfer in a simple, black-box manner. We detected information with uneven coverage across languages by controlling for presence and absence of Wikipedia articles in 12 languages. We generated knowledge-seeking questions in a source language, for which the answer appears in a relevant Wikipedia article and translated them to all other 11 languages, for which the respective Wikipedias lack equivalent articles. Assuming that Wikipedia reflects the prominent knowledge in the LLM’s training data, to solve ECLeKTic’s CBQA task the model is required to transfer knowledge between languages. Experimenting with 8 LLMs, we show that SOTA models struggle to effectively share knowledge across, languages even if they can predict the answer well for queries in the same language the knowledge was acquired in.
ECLeKTic: a Novel Challenge Set for
Evaluation of Cross-Lingual Knowledge Transfer
Omer Goldman∗γβ, Uri Shaham∗γ, Dan Malkinγ, Sivan Eigerγ, Avinatan Hassidimγ, Yossi Matiasγ, Joshua Maynezδ, Adi Mayrav Giladyγ, Jason Riesaδ, Shruti Rijhwaniδ, Laura Rimellδ, Idan Szpektorγ, Reut Tsarfatyγ, Matan Eyalγ βBar-Ilan University γGoogle Research, δGoogle DeepMind {ogoldman, urishaham, matane}@google.com
@@ 1 Introduction
Ideally, multilingual large language models (LLMs; Gemini Team, 2024; Llama Team, 2024; OpenAI, 2024, inter alia) should perform similarly and consistently in all languages they were trained on and specifically be as knowledgeable across these languages. In addition to making models more human-like, this would enable speakers of languages with smaller resource footprints on the Web to have equitable access to the world’s knowledge, and speakers of higher resource languages, to access a wider range of information. Alas, LLMs are known to be inconsistent across languages (e.g., Ohmer et al., 2023), and in particular, their performance on factual queries varies significantly depending on the language in which the model is queried Jiang et al. (2020); Kassner et al. (2021); Qi et al. (2023).
LLMs may learn similar knowledge in different languages when such documents are encountered during pretraining. But, since observing all of the world’s knowledge in all languages, e.g., via translation, is not feasible, a complementary requirement from LLMs may be to share acquired knowledge between languages, irrespective of the language it was acquired in. Models with such capability would either represent knowledge in a language-agnostic way, or implicitly translate knowledge at inference time from the language in which it was acquired to the target language. However, measuring this knowledge-sharing capability is not trivial. Methods suggested in the literature require either causal interventions that are far from perfect and limited in their transparency Ifergan et al. (2024); Wei et al. (2024) or careful dissection of the model’s inner states that may be noisy Chen et al. (2024); Zhao et al. (2024a).
@@
In this paper, we address the problem of empirically quantifying the cross-lingual knowledge-sharing ability of LLMs. Aiming for a simple, black-box evaluation, we introduce ECLeKTic to Evaluate Cross-Lingual Knowledge Transfer, by means of closed book QA. Consider, for example, a question asking who usually dubs characters played by Brad Pitt in German movies. Since in the German part of the internet it is well-known that the answer is Tobias Meister, a modern LLM is expected to answer this question easily when asked in German. However, to answer that in other languages, in which the internet contains little evidence of this fact, LLMs may struggle without being able to internally retrieve the German knowledge (see Figure 1).
To target facts like the one above, we constructed ECLeKTic, by targeting articles in Wikipedias in 12 languages that have no equivalent articles in the other languages, such as the article dedicated to Tobias Meister that exists only on the German Wikipedia. We generated fact-seeking question/answer pairs based on those articles using Gemini Gemini Team (2024) and translated them to the other tested languages. The entire generation and translation phases were manually verified by human annotators. The result is a set of questions and answers that relate to a fact that is well known only in one language (henceforth: the source language), but are contained in the dataset in all 12 languages (the target languages). Each question-answer pair is accompanied by the relevant Wikipedia context. Various LLMs were then asked these questions across all target languages in a closed-book setting and their predictions were judged by another model Chiang and Lee (2023); Zheng et al. (2023) in an open-book setting.
@@
Based on the predictions for each question in the 12 languages, we defined 2 metrics: overall success that reflects the model’s ability to solve ECLeKTic as a whole by transferring knowledge across languages, and transfer ability that only measures the model’s ability to transfer correct answers. We experimented with 8 top-performing models, both open-source and proprietary, to demonstrate that, across the board, ECLeKTic poses a significant challenge. The best performing model, Gemini 2.0 Pro, achieves overall success of 41.3% and manages to transfer only 65.0% of the facts it was able to retrieve in the the respective source language. Breaking down the results by source and target language, we show that shared script is a major factor in the ease of transfer, corroborating findings from previous works Qi et al. (2023); Ifergan et al. (2024). Finally, we tested models of various sizes, all from the Qwen 2.5 model series Qwen Team (2025), and found that bigger models are not able to transfer more knowledge in relative terms, i.e., in terms of transfer ability, although they are more successful in terms of overall success.
@@
All in all, the contribution of this paper is twofold. First, we introduce ECLeKTic, a novel benchmark for cross-lingual transfer evaluation, along with its construction and evaluation process. Second, we present a systematic evaluation of state-of-the-art models on ECLeKTic, showing lack of knowledge transfer across languages, leaving significant headroom for further research towards more capable and consistent multilingual models.
2 The Challenge of Cross-lingual Transfer Evaluation
With the rapid evolution of open-source and proprietary LLMs, we need robust black-box methods for evaluating and scrutinizing various model capabilities. In this work, we target the cross-lingual knowledge transfer abilities of multilingual LLMs. Specifically, we wish to assess whether consistent outputs for the same input in different languages occur due to genuine knowledge sharing — stemming from internal orthogonal treatment of knowledge and language — or due to incidental exposure and memorization of the same information in multiple languages during training.
To the best of our knowledge, prior work for directly assessing parametric knowledge-sharing and cross-lingual knowledge-transfer capabilities mainly looked at the internal mechanisms of open-source models. One line of work utilizes knowledge editing to determine whether the model’s representation of knowledge is causally linked across languages Ifergan et al. (2024); Wei et al. (2024). However, editing methods are far from perfect as the field is still evolving. Another approach uses direct observation of neuron activations in order to understand the extent to which actual transfer occurs Chen et al. (2024); Zhao et al. (2024a). One downside of this approach is that models’ inner states are not easily interpretable. Yet, more limiting is the fact that all the above methods are white-box inspection approaches that can only be applied to open-source models. These methods cannot assess modern SOTA proprietary models, for which only black-box analysis is possible.
In this work, we present ECLeKTic to black-box evaluate cross-lingual knowledge transfer using a well-established method of closed-book QA to query parametric knowledge AlKhamissi et al. (2022) by carefully selecting questions that would indicate genuine knowledge transfer.
3 ECLeKTic
In this section we introduce the ECLeKTic benchmark. We detail the construction of the dataset and then describe the evaluation procedure and metrics.
3.1 The ECLeKTic Dataset Construction
As ECLeKTic is a QA benchmark, we want to include only questions whose answers were exposed to the model in a single language during pre-training. Then, when we query such questions in other languages, an LLM could answer them from its parametric memory if it has a representation of that knowledge that is language agnostic.
@@ 3.1 The ECLeKTic Dataset Construction
To generate such question/answer pairs, we selected articles in Wikipedia that exist only in one of the 12 target languages (see list in Table 1). Concretely, we analyzed the July 2023 Wikipedia dump and for each language sampled 100 articles that contain at least 200 characters, had at least 100 views during 2023,222Statistics available on: https://stats.wikimedia.org/#/all-wikipedia-projects. and most importantly do not have equivalent articles in any of the other 11 languages. From each such article we extracted the 10 first sentences and based on them we instructed Gemini to generate a question and an answer.
@@
The context, the question, and the answer of all candidate generations were validated by human annotators. First, the annotators checked that the question is answerable in a closed book setting, i.e., it does not refer explicitly to the context or mention the answer. Second, they validated that the question is related to a fact that is particularly relevant to the language in question, e.g., it does not relate to a science or other general knowledge fact. Questions and answers that did not meet these criteria were discarded. Third, the annotators made sure that the question contains all the information needed to be answerable when translated. For example, a question in Hebrew relating to the TV series "Survivor" was disambiguated by the annotators to explicitly mention "the Israeli adaption of Survivor". Named entities were also clarified similarly, so a question referring to "Ambev" was modified refer to "the Brazilian brewing company, Ambev".
@@
Finally, each retained question, answer and context were automatically translated to the other 11 languages. The translations were verified by another set of human annotators and modified when needed. At this stage some examples were also discarded if they were proved to be untranslatable. For example, when a question explicitly refers to the meaning of a word in the source language. To overcome the difficulties in the verification of translation between non-trivial language pairs, this stage was done through English as a pivot language.
@@
The complete annotation process is depicted in Figure 2. All prompts that were used in the data creation are detailed in Appendix A. The statistics of the final benchmark are depicted in Table 1.
| Source Language | # Examples |
| English | 39 |
| French | 22 |
| German | 16 |
| Hebrew | 29 |
| Hindi | 64 |
| Indonesian | 31 |
| Italian | 29 |
| Japanese | 26 |
| Korean | 33 |
| Mandarin Chinese | 35 |
| Portuguese | 28 |
| Spanish | 32 |
| total generated | 384 |
| total translated | 4224 |
| total evaluated | 4608 |
3.2 ECLeKTic Metrics
@@
where is the prediction of model for question in language . is a 0/1 score that indicates whether questions that are expected to be correctly answered in language are indeed answered correctly both in this language as well as in another target language , capturing positive knowledge sharing between the two languages. Given a set of examples , the final metric overall simply averages across all question/answer pairs.
@@
Note that this metric does not explicitly reflect the number of questions answered correctly in their source language. As a result, it can be maximized even by weak models, for which is small, as long as they can answer the same questions in all languages.
To determine whether a model gives the correct answer to a specific question in a specific language we use an LLM as a judge Zheng et al. (2023). This is done in order to avoid the pitfalls of automatic metrics that are less correlated with human judgements Chen et al. (2019). The judge model, in our case Gemini 2.0 Flash, receives as input the question and the prediction of the tested model as well as the translated context in order to verify that the predicted answer is correct Zhou et al. (2025). The prompt used for judgement is in Appendix A.
3.3 Assumptions in ECLeKTic
First, we assume that all the tested models are exposed to the articles we extracted from Wikipedia. Moreover, we assume that the model was exposed to the information in those articles in their respective source languages multiple times during its training and the knowledge is therefore more accessible to it. In practice, Wikipedia itself is repeated in the pre-training data of most LLMs due to the quality of its texts Brown et al. (2020); Chowdhery et al. (2023), and we assume that it holds also for all multilingual models. Additionally, we assume that the existence of a Wikipedia article reflects a general interest of online speakers in the same topic so the information is also likely to be repeated in the same language outside of Wikipedia. This is even more straightforward given that we only targeted articles with significant yearly view count.
Conversely, we assume that the absence of an article from a certain Wikipedia reflects the lack of interest of online speakers in that topic. The information on that topic is therefore assumed to appear sparsely on the internet, if at all, and be far less accessible to the model in that language.
Together, these assumptions allow us to treat the (in)existence of a fact in Wikipedia as an approximate to the (in)exposure of a fact in a specific language to a multilingual model that was trained on the entire internet.
@@ 4.1 LLMs Struggle with Knowledge Transfer
| Model | Overall | Transfer |
| Gemini 2.0 Pro | 41.61.5 | 65.01.8 |
| GPT 4o | 38.81.4 | 67.01.8 |
| Gemini 2.0 Flash | 34.61.4 | 62.31.9 |
| Claude 3.5 Sonnet | 34.41.4 | 60.81.9 |
| Gemma 2 9B | 8.70.8 | 40.33.1 |
| Mistral Nemo | 7.10.8 | 38.93.4 |
| Qwen 2.5 7B | 2.80.5 | 23.53.7 |
| Olmo 2 7B | 1.60.3 | 17.23.7 |
To demonstrate ECLeKTic’s value in cross-lingual transfer evaluation, we measured the performance of several models. We included open models, namely the latest versions of Gemma,333https://huggingface.co/google/gemma-2-9b-it Mistral,444https://huggingface.co/mistralai/Mistral-Nemo-Instruct-2407 Qwen555https://huggingface.co/Qwen/Qwen2.5-7B-Instruct, and Olmo,666https://huggingface.co/allenai/OLMo-2-1124-7B-Instruct, all in sizes of between 7 and 9 billion parameters that underwent instruction tuning. In addition, the black-box nature of ECLeKTic allows the evaluation of closed models as well. We therefore included also GPT 4o,777We used version gpt-4o-2024-11-20. Claude 3.5 Sonnet,888Version claude-3-5-sonnet-20241022. and Gemini 2.0 in both Pro999Currently an experimental release. and Flash101010Version gemini-2.0-flash-001 versions.
All models were evaluated in a zero-shot setting and were prompted with the question on its own, i.e., without an explicit instruction. The results of can be found in Table 2. They show a clear gap in performance between the two groups, with proprietary models clearly outperforming the open ones by a very wide margin, probably due to their bigger size in terms of parameters (see also Section 4.3).
Gemini 2.0 Pro is the best performing model in solving the task of ECLeKTic, as defined in terms of overall success. It manages to answer correctly, in both source and target language, 41.3% of the examples. This score reveals that there is still a clear room for improvement in knowledge retrieval and transfer capabilities of all models. In terms of the transfer score, that is the portion of questions that were answered correctly in the target language out of those that got the right answer in their source languages, the performance in much better, with all closed models achieving more than 60%, but it is still far from perfect.
All in all, we conclude that ECLeKTic presents a serious challenge to modern LLMs despite their impressive abilities overall.

### v3

@@ Abstract
To achieve equitable performance across languages, large language models (LLMs) must be able to abstract knowledge beyond the language in which it was learnt. However, the current literature lacks reliable ways to measure LLMs’ capability of such cross-lingual knowledge transfer. To that end, we present ECLeKTic,111The dataset is available at https://www.kaggle.com/datasets/googleai/eclektic a multilingual closed-book QA dataset that Evaluates Cross-Lingual Knowledge Transfer in a simple, black-box manner. Concretely, we used the presence and absence of Wikipedia articles in 12 languages to detect pieces of information that were likely available during pre-training in one of the languages but not in the others. We curate ECLeKTic as a set of fact-seeking questions over this kind of information, in all the different languages. Therefore, in order to solve ECLeKTic the model is required to transfer knowledge between languages. We evaluated 8 LLMs and showed that current SOTA models struggle to effectively share knowledge across languages, even if they can predict the answer for questions in the language in which the knowledge was acquired.
@@ 1 Introduction
Multilingual large language models (LLMs; Gemini Team, 2024; Llama Team, 2024; OpenAI, 2024, inter alia) should perform similarly and consistently in all languages they were trained on, and crucially, possess an equivalent depth of knowledge across these languages. This would make models more similar to humans that decouple knowledge from language. But more importantly, it would make LLMs more efficient and provide equitable access to a wider range of the world’s knowledge to speakers of all languages.
However, measuring knowledge-sharing across languages is nontrivial. When a model provides consistent answers to the same question across different languages, it’s difficult to determine the underlying cause. Does this consistency stem from a single, language-independent representation of knowledge – or did the model simply learn from its training data the same fact in each language separately? Due to this ambiguity, previous works focused on cross-lingual consistency alone, measuring whether models answer the same across languages, without investigating the reason for that consistency (Jiang et al., 2020; Kassner et al., 2021; Ohmer et al., 2023; Qi et al., 2023). These works translated entire datasets, inherently testing facts that the model was likely exposed to in multiple languages during training, thus making it impossible to distinguish true knowledge sharing from parallel exposure.
In order to determine whether cross-lingual consistency in LLMs is incidental or principled we present ECLeKTic, to Evaluate Cross-Lingual Knowledge Transfer in a simple, black-box manner. We achieve that by building a closed-book Question Answering (QA) evaluation set that carefully targets specific facts that were likely acquired only in one language. Consider, for example, a question asking
@@
who usually dubs characters played by Brad Pitt in German movies. Since in the German part of the internet it is well-known that the answer is Tobias Meister, a modern LLM answers this question easily when asked in German. However, to answer that question in other languages, in which the internet contains little or no evidence of the underlying fact, LLMs that are not able to internally retrieve and share across languages that German knowledge may struggle (see Figure 1).
In this paper we present ECLeKTic to evaluate cross-lingual transfer by targeting such language-specific facts. We identified those facts using Wikipedia articles in 12 languages that lack equivalents in the other 11, such as the article on the aforementioned Tobias Meister, that exists only on the German Wikipedia. We annotated each such article with a fact-seeking question and answer, and translated them to all other tested languages in our benchmark. The entire generation and translation process was done by human annotators, based on LLM-generated suggestions. The result is a set of 4,224 question/answer pairs, each addressing a fact primarily known in one language (henceforth, the source language) but written in one of the other 11 language (the target languages). Additionally, each question-answer pair is accompanied by the Wikipedia paragraphs on which it was based, in its original (source) language.
We evaluated LLMs on our benchmark using a judge model (Chiang & Lee, 2023; Zheng et al., 2023), and aggregate example-level correctness into two metrics: overall success that reflects the ability to solve ECLeKTic and answer questions in their source and target languages, and transfer ability that only measures the ability to transfer correct answers.
@@
We experimented with 8 top-performing models, open-source and proprietary, and our results demonstrate that ECLeKTic poses a significant challenge across the board. The best performing model, Gemini 2.0 Pro, achieves overall success of 41.6% and manages to transfer only 65.0% of the facts it was able to retrieve in the respective source language. Breaking down the results by source and target language, we show that shared script is a major factor in the ease of transfer, corroborating findings from previous works (Malkin et al., 2022; Qi et al., 2023; Ifergan et al., 2024). Moreover, we showed, by examining models of different sizes from the Qwen 2.5 model series (Qwen Team, 2025), that that bigger models are not able to transfer more knowledge in relative terms, i.e., in terms of transfer ability, although they are more successful in terms of overall success.
@@
All in all, this work contributes a novel benchmark for cross-lingual transfer evaluation, along with its construction, metric definition, and evaluation process. ECLeKTic shows that there is a considerable headroom for further research into more capable and consistent multilingual models.
2 Constructing ECLeKTic
@@
ECLeKTic is a QA benchmark that includes questions of which answers were exposed to the model in a single language during pre-training. As a basis for such question/answer pairs, we selected articles in Wikipedia that only exist in one of the 12 target languages, and not in any of the others (English, French, German, Hebrew, Hindi, Indonesian, Italian, Japanese, Korean, Mandarin Chinese, Portuguese, and Spanish). Concretely, we analyzed the July 2023 dump of Wikipedia and for each language sampled 100 articles that contain at least 200 characters, had at least 100 views during 2023,222Statistics available on: https://stats.wikimedia.org/##/all-wikipedia-projects. and most importantly do not have equivalent articles in any of the other 11 languages.
@@
The annotation process was done in two stages. In the first, annotators were presented with Gemini-suggested a question-and-answer pair based on the 10 first sentences of each Wikipedia page. The annotators verified that the questions are based on the context, answerable from the context, and that the question is particularly relevant to the language in question, e.g., it does not relate to general common knowledge or scientific facts. The annotators also checked whether the question is answerable in a closed book setting, i.e., it does not refer explicitly to the context or mention the answer. Questions and answers that did not meet these criteria were discarded. Lastly, the annotators made sure that questions contain all the information needed to be answerable when translated. For example, a question in Hebrew relating to “the supreme court” was disambiguated to explicitly mention “the Israeli supreme court”. Named entities were also clarified similarly, so a question referring to “Ambev” was modified to refer to “the Brazilian brewing company, Ambev”.
@@
In the second stage, the questions, answers and their contexts were automatically translated to the other 11 languages by Gemini. Then they were given to another set of translation-expert annotators for verification and modification. Specifically, the annotators corrected the translation of uncommon named entities and verified that appropriate transliteration rules were applied. At this stage some examples were also discarded if they proved to be untranslatable. For example, when a question explicitly refers to the meaning of a word in the source language. To overcome the difficulties in translation between non-trivial language pairs, this stage was done through English as a pivot language. The annotators of this stage were also tasked with verifying the work of the previous annotators and discard any questions that do not comply with any of the aforementioned criteria. All in all, the human annotation effort amounted to 400 cumulative hours, performed by a team of 27 annotators.
@@
The complete annotation process is depicted in Figure 2. The prompts that were given to the model and the annotation guidelines that were given to the annotators in the data creation process are detailed in Appendices A and B.
2.1 Assumptions in ECLeKTic
First, we assume that all the tested models are exposed to the articles we extracted from Wikipedia. Moreover, we assume that models were exposed to the information in those articles in their respective source languages multiple times during their pretraining. This assumption is straightforward given that in practice Wikipedia itself is repeated in the pre-training data of most LLMs due to the quality of its texts (Brown et al., 2020; Chowdhery et al., 2023). Moreover, the existence of a Wikipedia article is likely to reflect a general interest of online speakers in the same topic so the information is also likely to be repeated in the same language outside of Wikipedia. This is even more straightforward given that we only targeted articles with significant yearly view count. Conversely, we assume that the absence of an article from a certain Wikipedia reflects the lack of interest of online speakers in that topic. The information on that topic is therefore assumed to appear sparsely on the internet, if at all.
Importantly, we assume a link between the abundance of information in a model’s training data and the accessibility of that information to the model from its weights. That is, we assume that without means of cross-lingual transfer it will be easier for a model to retrieve information in a certain language when it has seen this information more frequently in that language during pretraining.333See Verma et al. (2024) for a verification of a similar hypothesis regarding frequency and accessibility in computer vision.
Together, these assumptions allow us to treat the (non)existence of a fact in Wikipedia as an approximate to the (in)accessibility of that fact to a model trained on online data. In other words, we can treat questions about topics missing in Wikipedia in a certain language, as questions that require cross-lingual transfer to be answered.
2.2 The ECLeKTic Benchmark
The resulting benchmark includes 4608 question/answer pairs in 12 languages. 4224 of which are evaluated as target examples, i.e., examples in languages where the fact is not available to the model, and 384 are questions in their source language, used to evaluate the amount of information to be transferred. The abundance of each source language in ECLeKTic is in Figure 3.
Figure 4 details the distribution of questions across 10 domains detected automatically by Gemini.444We verified the domain classification by sampling 100 questions. All of them were classified correctly by Gemini. It reflects the various topics covered by ECLeKTic and verifies that our method of construction did not result in homogeneous questions. When breaking down the questions in ECLeKTic by domain separately for each source language, it is possible to see the we indeed captured the unique interests of speakers of each language. For example, "religion, philosophy, and mythology" is over-represented in the examples originated in Hebrew and Hindi, in accordance to the unique distribution of religions among their speakers, "arts and culture" is over represented in French and Italian, and "history" – in Chinese. The domain breakdown per-language is detailed in Appendix D.
3 ECLeKTic Metrics
@@
where is the prediction of model for question in language . is a binary score that indicates whether a question is correctly answered in its source language as well as in another target language , capturing positive knowledge sharing between the two languages. Given a set of examples , the final metric overall simply averages across all question/answer pairs.
@@
Note that this metric does not reflect the number of questions answered correctly in their source language. As a result, it can be maximized even by weak models, for which is small, as long as they answer the same questions in all languages.
To determine whether a model gives the correct answer to a specific question in a specific language we use an LLM as a judge (Zheng et al., 2023). Specifically, we let Gemini 2.0 Flash verify the answer’s correctness based on the question and the translated context (Zhou et al., 2025).555The prompt used for judgement is in Appendix B.
3.1 String Matching Evaluation
Evaluation by an LLM tends to be better than automatic metrics that are less correlated with human judgements (Chen et al., 2019). However, we introduce an automatic string matching metric to accompany the model-based metric.666The code for string-based evaluation of predictions can be found at https://github.com/omagolda/ECLeKTic-string-metrics This is done due to its ease of execution and to verify that the judge model does not introduce biases of its own.
In this metric we define the correctness of answer as word-level recall , i.e., the portion of words in the gold answer that are in the model’s prediction .777In Chinese and Japanese we counted every Chinese character as a word, for simplicity’s sake. Subsequently, the example-level success is then defined as
then remain the averaged example-level success over all examples. But as we can not say with this metric which examples are in , we amend to be a weighted average over the entire benchmark where the recall over the source examples serve as weights
@@ 4 Experiments
We used ECLeKTic to evaluate cross-lingual knowledge transfer in several leading models. We included open models, the latest instruction-tuned versions of Gemma,888https://huggingface.co/google/gemma-2-9b-it Mistral,999https://huggingface.co/mistralai/Mistral-Nemo-Instruct-2407 Qwen101010https://huggingface.co/Qwen/Qwen2.5-7B-Instruct, and
Olmo,111111https://huggingface.co/allenai/OLMo-2-1124-7B-Instruct all in sizes of between 7 and 9 billion parameters. In addition, the black-box nature of ECLeKTic allows the evaluation of closed models. We therefore included also GPT 4o,121212We used version gpt-4o-2024-11-20. Claude 3.5 Sonnet,131313Version claude-3-5-sonnet-20241022. and Gemini 2.0 in Pro141414Currently an experimental release. and Flash151515Version gemini-2.0-flash-001 versions.
All models were evaluated on the question on its own without an explicit instruction in the input.
| Model | LLM as a Judge | String Based | ||
| Overall | Transfer | Overall | Transfer | |
| Gemini 2.0 Pro | 41.61.5 | 65.01.8 | 41.41.5 | 61.21.7 |
| GPT 4o | 38.81.4 | 67.01.8 | 33.71.4 | 58.01.8 |
| Gemini 2.0 Flash | 34.61.4 | 62.31.9 | 32.11.4 | 56.01.8 |
| Claude 3.5 Sonnet | 34.41.4 | 60.81.9 | 24.91.3 | 49.11.9 |
| Gemma 2 9B | 8.70.8 | 40.33.1 | 9.60.9 | 37.42.4 |
| Mistral Nemo | 7.10.8 | 38.93.4 | 10.10.9 | 37.02.4 |
| Qwen 2.5 7B | 2.80.5 | 23.53.7 | 9.70.9 | 36.22.4 |
| Olmo 2 7B | 1.60.3 | 17.23.7 | 1.50.4 | 15.12.6 |
@@ 4.1 LLMs Struggle with Knowledge Transfer
The results can be found in Table 1. They show a clear gap in performance between the two groups, with proprietary models clearly outperforming the open ones by a very wide margin, probably due to their bigger size in terms of parameters (see also subsection 4.3).
Gemini 2.0 Pro is the best performing model, as defined in terms of overall success. It manages to answer correctly, in both source and target language, 41.6% of the examples. This score reveals that there is still a clear room for improvement in knowledge retrieval and transfer in all models. In terms of transfer score, the trends are pretty similar, with GPT 4o leading the pack. But even this more permissive metric shows that the model failed to transfer the knowledge of 37% of the questions correctly answered in their source language.
When comparing the metrics based on an LLM as a judge to the ones based on string matching, it is evident that the trends are largely similar: closed models are better than open ones and the ranking is between the models is similar. It is also shows that the Gemini instance used as a judge did not particularly favor the judged Gemini models. In fact, the opposite may be true, as the non-Gemini models were favored by the LLM judge compared to their string based evaluation.
All in all, we conclude that ECLeKTic presents a significant challenge to modern LLMs despite their impressive abilities overall.